This is a Biochemistry 200 level question, please answer neatly and as soon as possible. Thanks in advance, and I will like as soon as answer is posted!
The enzyme happyase catalyzes the following reaction: SAD — HAPPY An enzyme kinetics experiment gave the data shown in the taQUESTION 8 At what SAD concentration will the reaction rate equal 19.8 mM/s? Assume that the concentration of happyase is 25.
Answer #1

Following is the - complete Answer -&- Explanation: for the given Question, in.....typed format...

\RightarrowAnswer:

At reaction rate: Vo = 19.8 mM/sec, molar concentration of SAD, will be: [SAD] =  10.338 mM ( milli-mol/L )

\RightarrowExplanation:

Following is the complete Explanation, forr the above Answer....

  • Given:
  1. balanced chemical reaction:   SAD  \rightleftharpoons HAPPY
  2. Name of the substrate: SAD   
  3. Name of the product: HAPPY
  4. Name of the enzyme: happyase
  5. Enzyme concentration throughout the reaction:[Happyase] =  25.0 \mu M = 0.025 mM ( milli-mol/L )
  6. Vo =   Reaction Velocity [=] mM/s ( milli-mol/L .sec )
  7. [SAD] = Molar concentration of the substrate, in millimol/L ( i.e. mM )
  8. The values of [SAD], and the corresponding values of reaction velocity ( Vo ), is given in the question, as prodided below in the section: "Explanation"  
  • ​​​​​​​Step - 1:

​​​​​​​Following is the given information, in the form of a table, prepared by using MS Excel, presented here in...image format....

(SAD), IM (millimol/L) Vo, M/s (milli-mol/L. sec ) 3.7 6.7 9.25 11.385

Where:

  1. [SAD] = Molar concentration of the substrate, in millimol/L ( i.e. mM )
  2. Vo =   Reaction Velocity [=] mM/s ( milli-mol/L .sec ) ​​​​​​​
  • Step - 2:

​​​​​​​Since the plot, of Vo vs. [SAD], will not be Linear, we will have to plot the reciprocal values, i.e. the values of 1/Vo vs. 1/[SAD], which should be Linear. Therefore, we will have to prepare the following table, by modifyig the previous, given table,above as following, in....image format...

1/[SAD), mM^-1 ( L/millimol) 1/Vo, s/mM ( sec.1/milli-mol) 0.5 0.333333333 0.25 0.27027027 0.149253731 0.1081081081 0.0878348

  • Step - 3:

​​​​​​​Now, let's plot, the values of 1/Vo vs. 1/[SAD], according to the table, given above, using MS Excel, given here in...

....image format....

1/Vo vs. 1/[SAD] plot V = 0.243x + 0.0271 1/Vo, sec/mM 1/Vo vs. 1/(SAD) plot Linear (1/Vo vs. 1/(SAD) plot) 0.2 0.8 1.2 0.6 1

We get the following information, from the above plot...

  1. Equation of the obtained LinearTrendline: y = 0.243 x + 0.027   -------------------- Eq. (1)
  2. y - axis:   1/Vo   
  3. x - axis:   1/[SAD]
  • ​​​​​​​Step - 4:

​​​​​​​Therefore, the above Equation - (1), can be represented, as the following:

\Rightarrowi.e. 1/Vo = ( 0.243 ) x 1/[SAD] + 0.027   -------------------------------------- Eq. (2)

  • Step -5:

​​​​​​​Now, we are given:

\Rightarrow Reaction rate:   Vo = 19.8 mM/s  , i.e. under this condition, we will get the following:

\Rightarrow1/Vo = 1.0 / ( 19.8 mM/s ) =  0.050  ( approx. )

Plugging in the above obtained value in the given equation: Equation - (2), we will get the following:

\Rightarrow ( 0.050 ) = ( 0.243 ) x 1/[SAD] + 0.027

\Rightarrow1/[SAD] = [ (0.050 - 0.027 ) / ( 0.243 ) = 0.09672

Therefore, at that point in time, we will get the following value of [SAD].....

\Rightarrow[SAD] = 1.0 / ( 0.09672 ) =  10.338 mM  ( milli-mol/L )

  • Answer:

​​​​​​​Therefore, the molar concentration of SAD, will be the following, when the reaction velocity ( Vo ) , will be equal to 19.38 mM/s  

\Rightarrow[SAD] = 10.338 mM ( milli-mol/L )

-----------------------------------------------------------------------------------------------------------